# On the singular loci of higher secant varieties of Veronese embeddings

### v2

@@ 2.2. General secant fiber of a subsecant variety, entry locus, and its Veronese image
where $\pi_{x_{1}^{\prime}}:{\mathbb{P}}^{m}\dashrightarrow{\mathbb{P}}^{m-1}$ is the linear projection from $x_{1}^{\prime}$, and $v_{2}:{\mathbb{P}}^{m-1}\hookrightarrow{\mathbb{P}}^{\frac{(m+1)m}{2}-1}$ is the Veronese embedding of ${\mathbb{P}}^{m-1}$. Then $\dim(\tilde{\pi}(v_{2}({\mathbb{P}}^{m})))=\dim(v_{2}({\mathbb{P}}^{m-1}))=m-1$. If $k-1\leq\beta_{2}-2m$, then $\operatorname{codim}(\tilde{\pi}(v_{2}({\mathbb{P}}^{m})),{\mathbb{P}}^{\beta_{2}-m-1})-(k-1)\geq(\beta_{2}-2m)-(k-1)\geq 0$; hence the trisecant lemma implies $\tilde{\pi}(D)=\tilde{\pi}(v_{2}(x_{l}^{\prime}))$ for some $l=2,\dots,k$ as we discussed for (10).
In the above diagram, for any $y^{\prime}\in\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle\subset{\mathbb{P}}^{m}$ with $y^{\prime}\neq x_{1}^{\prime}$, we have $\tilde{\pi}(v_{2}(y^{\prime}))=v_{2}(\pi_{x_{1}^{\prime}}(y^{\prime}))=v_{2}(\pi_{x_{1}^{\prime}}(x_{l}^{\prime}))=\tilde{\pi}(v_{2}(x_{l}^{\prime}))$; indeed, $\tilde{\pi}^{-1}(\tilde{\pi}(v_{2}(x_{l}^{\prime})))=v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)$. Since $\tilde{\pi}(D)=\tilde{\pi}(v_{2}(x_{l}^{\prime}))$, we have $D\subset v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)$. ∎
Now we give a proof for Proposition 15. For simplicity, we set $i=1$. Then $x_{1}\in q_{1}(F)\subset Z=v_{d}({\mathbb{P}}^{m})$.
@@
(i) Set $s=\dim\sigma_{k}(Z)$. Then an irreducible component $F$ of $p^{-1}(a)$ is of dimension $\dim J-s=(km+k-1)-s$. From Lemma 13, we have $\dim q_{1}(F)=(km+k-1)-s$.
Let $q_{1}(F)^{\prime}\subset{\mathbb{P}}^{m}$ be the preimage of $q_{1}(F)\subset Z$ under $v_{d}:{\mathbb{P}}^{m}\simeq Z$, and let $A=q_{1}(F)^{\prime}\cup\left\{x_{1}^{\prime},\dots,x_{k}^{\prime}\right\}\subset{\mathbb{P}}^{m}$. Let $v_{d-1}:{\mathbb{P}}^{m}\rightarrow{\mathbb{P}}^{\beta_{d-1}}$ be the ($d-1$)-uple Veronese embedding. Then the ($k-1$)-plane $\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle\subset{\mathbb{P}}^{\beta_{d-1}}$ is contained in the linear variety
@@
and is of codimension $c=\dim\langle\,v_{d-1}(A)\,\rangle-(k-1)$. By Lemma 17 (i), $\beta_{d-1}-2m-k\geq 0$. So, by the generalized trisecant lemma, $v_{d-1}({\mathbb{P}}^{m})\cap\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle=\left\{v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\right\}$. In particular, $v_{d-1}(q_{1}(F)^{\prime})\cap\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle\subset\left\{v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\right\}$.
@@
Assume $d\geq 4$. Applying Lemma 19 (ii) to $R=\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime}),v_{d-1}(y^{\prime})\,\rangle\subset\langle\,v_{d-1}(A)\,\rangle$, we have
@@
In particular, $\dim_{v_{d-1}(x_{1}^{\prime})}(R\cap v_{d-1}(q_{1}(F)^{\prime}))=0$. On the other hand, regarding it as an intersection of two irreducible subvarieties in $\langle\,v_{d-1}(A)\,\rangle$, we deduce that every irreducible component of $R\cap v_{d-1}(q_{1}(F)^{\prime})$ is of dimension $\geq\dim(v_{d-1}(q_{1}(F)^{\prime}))-(c-1)$. Hence $\dim(\langle\,v_{d-1}(A)\,\rangle)\geq k+\dim(v_{d-1}(q_{1}(F)^{\prime}))=k+(km+k-1)-s$.
@@
is a linear subvariety of dimension $\leq k$. Since $v_{2}(q_{1}(F)^{\prime})\cap\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime})\,\rangle$ is empty or is a set of points, the intersection $v_{2}(q_{1}(F)^{\prime})\cap\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle$ is of dimension $\leq 1$. On the other hand, since $m\neq 2$ by Remark 18 (a), we have $\dim q_{1}(F)^{\prime}=(km+k-1)-\beta\geq 2$ as in Remark 20 below. For the union
@@
Let $R=\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(y^{\prime})\,\rangle\subset\langle\,v_{2}(A)\,\rangle$ and suppose that $\dim_{v_{2}(x_{1}^{\prime})}(R\cap v_{2}(q_{1}(F)^{\prime}))>0$, that is to say, there is a curve $C\subset R\cap v_{2}(q_{1}(F)^{\prime})$ containing ${v_{2}(x_{1}^{\prime})}$. Taking $D=C$ and applying Lemma 19 (iii), we have $C=v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)$ for some $l>1$. If $\dim\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle=k$, then $R=\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle$, contradicting to the definition of $W$ and our choice of $y^{\prime}$. Else if, $\dim\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle=k-1$, then $C=v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\subset v_{2}(q_{1}(F)^{\prime})\cap\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime})\,\rangle$, also contradicting that the intersection is of dimension at most 00.
Hence $\dim_{v_{2}(x_{1}^{\prime})}(R\cap v_{2}(q_{1}(F)^{\prime}))=0$. Then, in the same way as above, we have $\dim(\langle\,v_{2}(A)\,\rangle)\geq k+(km+k-1)-s$.
(ii) In the case when $(k-1)m+k=\binom{m+d}{m}$, we have $km+k-1-s\geq m$. It follows from Lemma 13 and ${\mathbb{P}}^{m}\simeq Z$ that $q_{i}(F)=Z$. From Lemma 17 (ii), if $(d,m)\neq(3,2)$, then ${\mathbb{P}}^{\beta_{d-1}}=\langle\,v_{d-1}({\mathbb{P}}^{m})\,\rangle$ is of dimension $\geq k+m$. ∎
Remark 20.
For $d=3$, $m\neq 2$, if $\mu:=\frac{\binom{m+d}{m}}{m+1}\notin{\mathbb{N}}$ and $k=\lceil\mu\rceil$ (as under the conditions of Proposition 15 and Remark 16 (a)), then we have $(km+k-1)-\beta_{d}\neq 1$. The reason is as follows. First, we may write $\mu={(m+3)(m+2)}/6=M/3$ for some $M\in{\mathbb{N}}$ since $(m+3)(m+2)$ is divided by $2$. In addition, dividing $M$ by $3$ with remainder, we have $M=3Q+R$ for a quotient $Q$ and a remainder $R$. Since $\mu=M/3\notin{\mathbb{N}}$, $R$ must be $1$ or $2$. In this setting, $k=\lceil M/3\rceil=Q+1$. It follows that $(km+k-1)-\beta_{d}$ is equal to
$(km+k)-\tbinom{m+3}{3}=(m+1)\left({k-\tfrac{(m+3)(m+2)}{6}}\right)=(m+1)\left((Q+1)-(Q+\tfrac{R}{3})\right)=(m+1)\cdot\tfrac{3-R}{3}~{}.$
We end this subsection by making one more important remark on the case of secant defective $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$’s, which do not satisfy the condition $\beta<km+k-1$ in Proposition 15.
@@
For four defective cases $(k,d,m)=(5,4,2)$,$(9,4,3)$,$(14,4,4)$ and $(7,3,4)$, similarly as in Proposition 15, we can have an estimation
$\dim\langle\,v_{d-1}(A)\,\rangle\geq k+\delta~{},$
where $A=v_{d}^{-1}(q_{1}(p^{-1}(a)))\subset{\mathbb{P}}^{m}$, the preimage of the entry locus of $a$, and $\delta$ is the secant defect of $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$; here $A$ is $\delta$-equidimensional, the $k$ general points $x_{1}^{\prime},\dots,x_{k}^{\prime}\in{\mathbb{P}}^{m}$ are contained in $A$, and it’s well known that $\delta=2$ when $(k,d,m)=(9,4,3)$ and $\delta=1$ in all the other defective cases.
For 3 cases $(k,d,m)=(5,4,2),(9,4,3)$, and $(14,4,4)$, we see that $\beta_{d-1}-2m\geq k$ and $v_{d-1}(A)\cap\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle\subset\left\{v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\right\}$ by the trisecant lemma so that we may take a point $y^{\prime}\in A$ such that $v_{d-1}(y^{\prime})\notin\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle$. By Lemma 19 (ii), we get $\dim_{v_{d-1}(x_{1}^{\prime})}(R\cap v_{d-1}({\mathbb{P}}^{m}))=0$ where $R=\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime}),v_{d-1}(y^{\prime})\,\rangle$. Thus, by the same intersection argument in $\langle\,v_{d-1}(A)\,\rangle$ (as Proposition 15 (i)), we derive the estimation
@@
For the remaining case $(k,d,m)=(7,3,4)$, it holds $\beta_{d-1}-2m=k-1$, and we still can claim that $\dim(\langle\,v_{2}(A)\,\rangle)\geq k+\delta=7+1=8$ as follows. For the $6$-dimensional subspace $M:=\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{7}^{\prime})\,\rangle\subset\langle\,v_{2}(A)\,\rangle$, the trisecant lemma implies $M\cap v_{2}(A)\subset M\cap v_{2}({\mathbb{P}}^{4})=\left\{v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{7}^{\prime})\right\}$, 00-dimensional intersection. Then $\dim(\langle\,v_{2}(A)\,\rangle)\geq 7$ (otherwise, we get $M=\langle\,v_{2}(A)\,\rangle$ so that $M\cap v_{2}(A)=v_{2}(A)$, a contradiction). Suppose that $\dim(\langle\,v_{2}(A)\,\rangle)=7$, and set $R=\langle\,v_{2}(A)\,\rangle$. We take the irreducible components of $1$-equidimensional closed set $A$ as $A=\bigcup_{j=1}^{s}A_{j}$. Note that $v_{2}(A_{j})\subset R\cap v_{2}({\mathbb{P}}^{m})$. Since $x_{1}^{\prime}\in A$, there is a curve $A_{j_{0}}$ containing $x_{1}^{\prime}$. Taking $C=v_{2}(A_{j_{0}})$ and applying Lemma 19 (iii), for any $j$ with $1\leq j\leq s$, we have $v_{2}(A_{j})=v_{2}(\langle x_{1}^{\prime},x_{l_{j}}^{\prime}\rangle)$ for some $l_{j}=2,\dots,k$. It is equivalent to $A_{j}=\langle x_{1}^{\prime},x_{l_{j}}^{\prime}\rangle$, a line in ${\mathbb{P}}^{4}$; in particular, $x_{1}^{\prime}\in A_{j}$. In the same way, $A_{j}$ must contain $x_{1}^{\prime},\dots,x_{7}^{\prime}$. But this is a contradiction, because these points are chosen as general $7$ points in ${\mathbb{P}}^{4}$. Thus, $\dim(\langle\,v_{2}(A)\,\rangle)\geq 8$.

### v3

@@ 2.2. General secant fiber of a subsecant variety, entry locus, and its Veronese image
where $\pi_{x_{1}^{\prime}}:{\mathbb{P}}^{m}\dashrightarrow{\mathbb{P}}^{m-1}$ is the linear projection from $x_{1}^{\prime}$, and $v_{2}:{\mathbb{P}}^{m-1}\hookrightarrow{\mathbb{P}}^{\frac{(m+1)m}{2}-1}$ is the Veronese embedding of ${\mathbb{P}}^{m-1}$. Then
$\dim(\tilde{\pi}(v_{2}({\mathbb{P}}^{m})))=\dim(v_{2}({\mathbb{P}}^{m-1}))=m-1.$
If $k-1\leq\beta_{2}-2m$, then
$\operatorname{codim}(\tilde{\pi}(v_{2}({\mathbb{P}}^{m})),{\mathbb{P}}^{\beta_{2}-m-1})-(k-1)\geq(\beta_{2}-2m)-(k-1)\geq 0;$
hence the trisecant lemma implies $\tilde{\pi}(D)=\tilde{\pi}(v_{2}(x_{l}^{\prime}))$ for some $l=2,\dots,k$ as we discussed for (15).
In the diagram above, for any $y^{\prime}\in\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle\subset{\mathbb{P}}^{m}$ with $y^{\prime}\neq x_{1}^{\prime}$, we have
$\tilde{\pi}(v_{2}(y^{\prime}))=v_{2}(\pi_{x_{1}^{\prime}}(y^{\prime}))=v_{2}(\pi_{x_{1}^{\prime}}(x_{l}^{\prime}))=\tilde{\pi}(v_{2}(x_{l}^{\prime}));$
indeed, $\tilde{\pi}^{-1}(\tilde{\pi}(v_{2}(x_{l}^{\prime})))=v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)$. Since $\tilde{\pi}(D)=\tilde{\pi}(v_{2}(x_{l}^{\prime}))$, we have $D\subset v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)$. ∎
We give one calculation before proving Proposition 15.
Remark 20.
For $d=3$ and $m\neq 2$, if $\mu_{0}=\frac{\binom{m+d}{m}}{m+1}\notin{\mathbb{N}}$ and $k=\lceil\mu_{0}\rceil$ (as under the conditions of Proposition 15 and Remark 16 (a)), then we have $(km+k-1)-\beta_{d}\neq 1$. The reason is as follows. First, we may write $\mu_{0}={(m+3)(m+2)}/6=M/3$ for some $M\in{\mathbb{N}}$ since $(m+3)(m+2)$ is divisible by $2$. In addition, dividing $M$ by $3$ with remainder, we have $M=3Q+R$ for a quotient $Q$ and a remainder $R$. Since $\mu_{0}=M/3\notin{\mathbb{N}}$, $R$ must be $1$ or $2$. In this setting, $k=\lceil M/3\rceil=Q+1$. It follows that $(km+k-1)-\beta_{d}$ is equal to
$(km+k)-\tbinom{m+3}{3}=(m+1)\left({k-\tfrac{(m+3)(m+2)}{6}}\right)=(m+1)\left((Q+1)-(Q+\tfrac{R}{3})\right)=(m+1)\cdot\tfrac{3-R}{3}.$
@@
(i) For simplicity, we set $i=1$; then $x_{1}\in q_{1}(F)\subset Z=v_{d}({\mathbb{P}}^{m})$. For $s=\dim\sigma_{k}(Z)$, an irreducible component $F$ of $p^{-1}(a)$ is of dimension $\dim J-s=(km+k-1)-s$. From Lemma 13, we have $\dim q_{1}(F)=(km+k-1)-s$.
Let $q_{1}(F)^{\prime}\subset{\mathbb{P}}^{m}$ be the preimage of $q_{1}(F)\subset Z$ under $v_{d}:{\mathbb{P}}^{m}\simeq Z$, and let
$A=q_{1}(F)^{\prime}\cup\left\{x_{1}^{\prime},\dots,x_{k}^{\prime}\right\}\subset{\mathbb{P}}^{m}.$
Let $v_{d-1}:{\mathbb{P}}^{m}\rightarrow{\mathbb{P}}^{\beta_{d-1}}$ be the ($d-1$)-uple Veronese embedding. Then the ($k-1$)-plane
$\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle\subset{\mathbb{P}}^{\beta_{d-1}}$
is contained in the linear variety
@@
and is of codimension $c=\dim\langle\,v_{d-1}(A)\,\rangle-(k-1)$. By Lemma 17 (i), $\beta_{d-1}-2m-k\geq 0$. So, by the generalized trisecant lemma,
$v_{d-1}({\mathbb{P}}^{m})\cap\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle=\left\{v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\right\}.$
In particular,
$v_{d-1}(q_{1}(F)^{\prime})\cap\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle\subset\left\{v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\right\}.$
@@
Assume $d\geq 4$. Applying Lemma 19 (ii) to
$R=\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime}),v_{d-1}(y^{\prime})\,\rangle\subset\langle\,v_{d-1}(A)\,\rangle,$
we have
@@
In particular, $\dim_{v_{d-1}(x_{1}^{\prime})}(R\cap v_{d-1}(q_{1}(F)^{\prime}))=0$. Regarding it as an intersection of two irreducible subvarieties in $\langle\,v_{d-1}(A)\,\rangle$, we deduce that every irreducible component of $R\cap v_{d-1}(q_{1}(F)^{\prime})$ is of dimension $\geq\dim(v_{d-1}(q_{1}(F)^{\prime}))-(c-1)$. Hence
$\dim(\langle\,v_{d-1}(A)\,\rangle)\geq k+\dim(v_{d-1}(q_{1}(F)^{\prime}))=k+(km+k-1)-s.$
@@
is a linear subvariety of dimension $\leq k$. Since $v_{2}(q_{1}(F)^{\prime})\cap\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime})\,\rangle$ is empty or is a set of points, the intersection
$v_{2}(q_{1}(F)^{\prime})\cap\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle$
is of dimension $\leq 1$. On the other hand, since $m\neq 2$ by Remark 18 (a), we have
$\dim q_{1}(F)^{\prime}=(km+k-1)-\beta\geq 2$
as in Remark 20. For the union
@@
Let $R=\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(y^{\prime})\,\rangle\subset\langle\,v_{2}(A)\,\rangle$ and suppose that $\dim_{v_{2}(x_{1}^{\prime})}(R\cap v_{2}(q_{1}(F)^{\prime}))>0$, that is to say, there is a curve $C\subset R\cap v_{2}(q_{1}(F)^{\prime})$ containing ${v_{2}(x_{1}^{\prime})}$. Taking $D=C$ and applying Lemma 19 (iii), we have $C=v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)$ for some $l>1$. If
$\dim\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle=k,$
then $R=\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle$, contradicting to the definition of $W$ and our choice of $y^{\prime}$. Else if
$\dim\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime}),v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\,\rangle=k-1,$
then $C=v_{2}(\langle\,x_{1}^{\prime},x_{l}^{\prime}\,\rangle)\subset v_{2}(q_{1}(F)^{\prime})\cap\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{k}^{\prime})\,\rangle$, also contradicting that the intersection is of dimension at most 00.
Hence $\dim_{v_{2}(x_{1}^{\prime})}(R\cap v_{2}(q_{1}(F)^{\prime}))=0$. Then, in the same way as above, we have
$\dim(\langle\,v_{2}(A)\,\rangle)\geq k+(km+k-1)-s.$
(ii) In the case when
$(k-1)m+k=\tbinom{m+d}{m},$
we have $km+k-1-s\geq m$. It follows from Lemma 13 and ${\mathbb{P}}^{m}\simeq Z$ that $q_{i}(F)=Z$. From Lemma 17 (ii), if $(d,m)\neq(3,2)$, then ${\mathbb{P}}^{\beta_{d-1}}=\langle\,v_{d-1}({\mathbb{P}}^{m})\,\rangle$ is of dimension $\geq k+m$. ∎
We end this subsection by making one more important remark on the case when $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subsetneq{\mathbb{P}}^{\beta_{d}}$ is secant defective, which will be used in the proof of of Part (ii) of Theorem 3.
@@
For four defective cases
$(k,d,m)=(7,3,4),(5,4,2),(9,4,3),(14,4,4),$
similarly as in Proposition 15, we can have an estimation
$\dim\langle\,v_{d-1}(A)\,\rangle\geq k+\delta,$
where $A=v_{d}^{-1}(q_{1}(p^{-1}(a)))\subset{\mathbb{P}}^{m}$, the preimage of the entry locus of $a$, and $\delta$ is the secant defect of $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$; here $A$ is $\delta$-equidimensional, the $k$ general points $x_{1}^{\prime},\dots,x_{k}^{\prime}\in{\mathbb{P}}^{m}$ are contained in $A$, and it is well known that $\delta=2$ when $(k,d,m)=(9,4,3)$ and $\delta=1$ in all the other defective cases.
For three cases
$(k,d,m)=(5,4,2),(9,4,3),(14,4,4),$
we see that $\beta_{d-1}-2m\geq k$ and
$v_{d-1}(A)\cap\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle\subset\left\{v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\right\}$
by the trisecant lemma so that we may take $y^{\prime}\in A$ such that $v_{d-1}(y^{\prime})\notin\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime})\,\rangle$. By Lemma 19 (ii), we get
$\dim_{v_{d-1}(x_{1}^{\prime})}(R\cap v_{d-1}({\mathbb{P}}^{m}))=0,$
where $R=\langle\,v_{d-1}(x_{1}^{\prime}),\dots,v_{d-1}(x_{k}^{\prime}),v_{d-1}(y^{\prime})\,\rangle$. Thus, by the intersection argument in $\langle\,v_{d-1}(A)\,\rangle$ (similar to Proposition 15 (i)), we derive the estimation
@@
For the remaining case $(k,d,m)=(7,3,4)$, it holds $\beta_{d-1}-2m=k-1$, and we still can claim that
$\dim(\langle\,v_{2}(A)\,\rangle)\geq k+\delta=7+1=8$
as follows. For the $6$-dimensional subspace $M=\langle\,v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{7}^{\prime})\,\rangle\subset\langle\,v_{2}(A)\,\rangle$, the trisecant lemma implies
$M\cap v_{2}(A)\subset M\cap v_{2}({\mathbb{P}}^{4})=\left\{v_{2}(x_{1}^{\prime}),\dots,v_{2}(x_{7}^{\prime})\right\},$
00-dimensional intersection. Then $\dim(\langle\,v_{2}(A)\,\rangle)\geq 7$ (otherwise, we get $M=\langle\,v_{2}(A)\,\rangle$ so that $M\cap v_{2}(A)=v_{2}(A)$, a contradiction). Suppose that
$\dim(\langle\,v_{2}(A)\,\rangle)=7,$
and set $R=\langle\,v_{2}(A)\,\rangle$. We take the irreducible components of the $1$-equidimensional closed set $A$ as $A=\bigcup_{j=1}^{s}A_{j}$. Note that $v_{2}(A_{j})\subset R\cap v_{2}({\mathbb{P}}^{m})$. Since $x_{1}^{\prime}\in A$, there is a curve $A_{j_{0}}$ containing $x_{1}^{\prime}$. Taking $C=v_{2}(A_{j_{0}})$ and applying Lemma 19 (iii), for any $j$ with $1\leq j\leq s$, we have $v_{2}(A_{j})=v_{2}(\langle x_{1}^{\prime},x_{l_{j}}^{\prime}\rangle)$ for some $l_{j}=2,\dots,k$. It is equivalent to $A_{j}=\langle x_{1}^{\prime},x_{l_{j}}^{\prime}\rangle$, a line in ${\mathbb{P}}^{4}$; in particular, $x_{1}^{\prime}\in A_{j}$. In the same way, $A_{j}$ must contain $x_{1}^{\prime},\dots,x_{7}^{\prime}$. But this is a contradiction, because these points are chosen as general $7$ points in ${\mathbb{P}}^{4}$. Thus, $\dim(\langle\,v_{2}(A)\,\rangle)\geq 8$.